# Systematics of neutron emission

### v2

@@ 4 Conclusions
Concluding, we analyzed the available experimental data concerning the neutron emission half life between ground states for light nuclei with $A\leq$ 28. In the case of charged particles it is well known the correlation between the logarithm of the half life and Coulomb parameter $\chi\sim ZQ^{-1/2}$, characterizing the irregular Coulomb wave. For neutral particles this function is replaced by Neumann spherical wave. It turns out that the half lives of light neutron emitters are rather well fitted by the spherical Neumann wave squared, depending on angular momentum estimated at the reduced nuclear radius $\rho\sim A^{1/3}Q^{1/2}$ corresponding to the experimental Q-value. Moreover, as in the case of the proton emission, the reduced half-lives estimated by excluding the centrifugal factor are proportional to the monopole spherical Neumann function squared. Therefore the well known dependence in the emission of charged particles $\log_{10}T_{/12}\sim Q^{-1/2}$ (Geiger-Nuttall law) is replaced in the emission of neutral particles by $\log_{10}T_{/12}\sim Q^{1/2}$. This is a remarkable feature taking into accound the relative short half lives of these neutron rich nuclei. It turns out that the estimated influence of the neutron formation probability is about a factor of three, while a difference by one unit in the angular momentum leads to an effect about two orders of magnitude in the half life. Therefore the investigation of the neutron emission can be a powerfull tool to assign the angular momentum of the unbound neutron.

### v3

@@ 4 Conclusions
Concluding, we analyzed the available experimental data concerning the neutron emission half life between ground states for light nuclei with $A\leq$ 28. In the case of charged particles it is well known the correlation between the logarithm of the half life and Coulomb parameter $\chi\sim ZQ^{-1/2}$, characterizing the outgoing Coulomb-Hankel wave. For neutral particles this function is replaced by the spherical Hankel wave. It turns out that the half lives of light neutron emitters are rather well fitted by the modulus squared of the spherical Hankel wave, estimated at the reduced nuclear radius $\rho\sim A^{1/3}Q^{1/2}$ corresponding to the experimental Q-value. Moreover, as in the case of the proton emission, the reduced half lives estimated by excluding the centrifugal factor are proportional to the modulus squared of the monopole spherical Hankel function. Therefore the well known dependence in the emission of charged particles $\log_{10}T\sim ZQ^{-1/2}$ (Geiger-Nuttall law) is replaced in the emission of neutral particles by $T\sim\rho^{-2}\sim A^{-2/3}Q^{-1}$ for monopole transitions. This is a remarkable feature taking into accound the relative short half lives of these neutron rich nuclei. It turns out that the estimated influence of the neutron formation probability is about a factor of three, while a difference by one unit in the angular momentum leads to an effect about two orders of magnitude in the half life. Therefore the investigation of the neutron emission can be a powerfull tool to assign the angular momentum of the unbound neutron.